# Evaluating The Explainability of State-of-the-Art Machine Learning-based Online Network Intrusion Detection Systems

### v2

@@ VI Discussion and Future Work
From the experimental results obtained in the previous section, we have seen that some ML-based NIDS models (Kitsune, ENIDrift) can be better interpreted in terms of decision trees than other models (HorusEye, HELAD). This might be attributed to the complexity of the NIDS model architecture and the properties of the model components. For example, HorusEye NIDS fidelity values are in the negative. Though HorusEye architecture includes only a single autoencoder in the second stage (which processes control plane traffic) compared to an ensemble of AEs in Kitsune or ENIDrift, that single AE is modified from a conventional AE to keep it lightweight, reduce the computational complexity, reduce the false-positive rate and improve throughput. It uses an asymmetric AE with separable convolution, dilation convolution and model quantization. As it is, AEs are limited in their interpretability as a result of their non-linear nature. Due to the complex architecture of the modified AE deployed in HorusEye, TRUSTEE might not be able to build a DT which can imitate the decisions of the black-box modified AE within an acceptable error rate. Therefore, to build ML-based NIDSs with better “explainability”, we recommend using less complex ML architectures and conventional components. However, there will be a trade-off between “explainability” and model performance (e.g., false-positive rate) and NIDS designers need to achieve a balance between the two.
@@
We have also seen from the results that TRUSTEE and SHAP explanations are in conflict for 2 out of 3 ML-based NIDSs considered in our work. Their explanations are in agreement for HorusEye NIDS but in conflict for Kitsune and ENIDrift NIDSs. This is probably due to the fact that SHAP value for each feature is calculated for a subset of data samples and reflects the importance of that feature in the decision made by the black-box model for a majority of the data samples in that subset. On the other hand, TRUSTEE acts on the complete input dataset and identifies the top features contributing the most to forming the decision boundary separating all benign and malicious data samples which are part of the input dataset. Another reason could be the different optimization criteria used by TRUSTEE and SHAP while building explanations. TRUSTEE applies a teacher-student dynamic derived from imitation learning that uses the black-box ML model as an oracle in conjunction with a carefully curated dataset to guide the training of a surrogate “white-box” model in the form of a DT that imitates the black-box’s decisions. On the other hand, SHAP uses the Shapley kernel method to approximate the black-box model-interpretable model fidelity function which recovers Shapley values. The specific data subset used to conduct the TRUSTEE-SHAP comparison could also play a role in the conflict. To reduce the conflict between TRUSTEE and SHAP explanations at the subset-level, we need to further investigate its underlying cause in the future.
@@
Finally, it can be inferred from the results that some ML-based NIDSs (Kitsune) are more vulnerable to o.o.d. data samples than others (HorusEye, HELAD) while for the rest (ENIDrift), such vulnerability remains inconclusive. One of the possible underlying reasons is that sometimes NIDSs are trained with training data containing higher than real-world attack traffic volume. To reduce vulnerability to o.o.d samples, we recommend that NIDS designers use training data which reflects the attack traffic volume in real-world deployments. Additionally, the NIDS architecture may also have a role to play in its vulnerability to o.o.d. samples. An AE trained on benign traffic may be vulnerable to o.o.d. attack traffic samples as the threshold used for output loss function (RMSE) is determined based on “in-distribution” attack traffic present in the training data. Since, Kitsune consists of an ensemble of AEs, the vulnerability effect is amplified. HorusEye consists of a single modified AE and therefore, its vulnerability to o.o.d. samples is not as pronounced as Kitsune. ENIDrift uses a different method of feature extraction (vector embedding) and updates the weights assigned to AEs in the ensemble based on change in attack traffic distribution which might keep it from producing a discernable pattern of dependence on attack traffic volume compared to Kitsune. We have observed earlier that HELAD NIDS anomaly scores are not vulnerable to o.o.d. samples to a significant extent. This may be attributed to HELAD’s architecture which combines an AE with an LSTM network. Since the final anomaly score is a weighted exponential aggregation of AE’s actual RMSE score and LSTM’s predicted RMSE scores for future samples (using historical samples). Even if the AE is vulnerable to o.o.d. samples, the LSTM’s predicted RMSE scores might be much larger compared to to the AE’s RMSE score (the exponential operation makes the value even larger), thus affecting the final anomaly score to a greater extent and reducing HELAD’s vulnerability to o.o.d. samples.
@@ VII Conclusion
We have analyzed four state-of-the-art ML-based online NIDS models (Kitsune, HorsusEye, ENIDRift, HELAD) using explainable AI techniques such as TRUSTEE and SHAP. Using both global explanations and local explanations for the blackbox models’ decisions, we have presented the most prominent features used by each NIDS model considered. We have compared the explanations generated across xAI methods for a given NIDS model (using our devised data subset-level explanation approach) as well as the explanations generated across the NIDS models for a given xAI method (using fidelity and other standard evaluation metrics). Finally, we evaluate the vulnerability of each NIDS model to inductive bias (o.o.d data samples). The results show that: (1) some ML-based NIDS models can be better interpreted in terms of decision trees than other models, (2) TRUSTEE and SHAP explanations are in conflict for most of the NIDS models considered in this work and (3) some NIDS models are more vulnerable to o.o.d samples than other models.
@@ Appendix A More Experimental Results
In this section, we present explanations of the four state-of-the-art ML-based online NIDSs (Kitsune, HorusEeye, ENIDrift, HELAD) generated using xAI methods such as TRUSTEE and kernel SHAP where the NIDSs were trained/tested with the CICIDS-2017 dataset.
@@ A-A Case Study- Kitsune $\times$ CICIDS-2017
TRUSTEE Analysis: The results of TRUSTEE analysis for Kitsune are shown in Table VI. Using TRUSTEE with $30\%$ of the original CICIDS-2017 dataset samples and no pruning results in the DT explanation which achieves 0.843 fidelity compared to Kitsune. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=3$) results in a DT explanation that is shown in Fig. 13 and achieves 0.919 fidelity compared to Kitsune.
@@
TABLE VI: TRUSTEE analysis results for Kitsune CICIDS-2017 model
@@
Detection of Inductive Bias: The CICIDS-2017 dataset does not provide a separate dataset for botnet traffic trace which could be used for comparison with the Mirai trace in terms of the attack traffic volume. However, using a similar approach for tampering as in the case of Mirai trace, we evaluated the following threshold ranges (in packets per second) for ARP requests originating from infected hosts : (1) 10-50, (2) 30-70 and (3) 50-90. The original trace as well as the tampered traces are shown in Fig. 15. It can be observed that: (1) there is no significant difference between the benign and attack traffic volumes for the original trace and (2) tampering had negligible effect on the difference between benign and attack traffic in terms of number of packets per second. Thus, unlike the Mirai trace, there is no significant difference between the benign and attack traffic volumes which can be learnt by the NIDS model and leave it vulnerable to a change in attack traffic distribution. Thus, we do not investigate the presence of inductive bias (vulnerability to o.o.d samples) any further for NIDS models trained with the CICIDS-2017 dataset.
@@ A-B Case Study- HorusEye $\times$ CICIDS-2017
TRUSTEE Analysis: The results of TRUSTEE analysis for HorusEye are shown in Table VII. Using TRUSTEE with $30\%$ of the CICIDS-2017 dataset samples with no pruning results in a DT explanation which achieves negative fidelity ($\sim-0.956$) compared to original HorusEye model. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation which achieves negative fidelity ($\sim-0.018$) compared to original HorusEye model but higher fidelity compared to no pruning. Negative fidelity values mean that the DT explanation fits HorusEye predictions worse than the mean of the predictions.
@@
TABLE VII: TRUSTEE analysis results for HorusEye CICIDS-2017 model
@@ A-C Case Study- ENIDrift $\times$ CICIDS-2017
TRUSTEE Analysis: The results of TRUSTEE analysis for ENIDrift are shown in Table VIII. Using TRUSTEE with $30\%$ of the CICIDS-2017 dataset samples and no pruning results in the DT explanation which achieves 0.929 fidelity compared to ENIDrift. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation (shown in Fig. 18) which achieves 0.53 fidelity compared to ENIDrift. This is expected since pruning the branches of a DT reduces its fidelity. The most prominent features ENIDrift uses to determine anomaly are: $f38$, $f81$, $f158$.
@@
TABLE VIII: TRUSTEE analysis results for ENIDrift CICIDS-2017 model
@@ A-D Case Study- HELAD $\times$ CICIDS-2017
TRUSTEE Analysis: The results of TRUSTEE analysis for HELAD are shown in Table IX. Using TRUSTEE with $30\%$ of the reduced CICIDS-2017 dataset samples and no pruning results in the DT explanation which achieves 0.0 fidelity compared to HELAD. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation (shown in Fig. 20) which achieves 0.0 fidelity compared to HELAD.
@@
TABLE IX: TRUSTEE analysis results for HELAD NIDS

### v3

@@ VI Discussion and Future Work
From the experimental results obtained in the previous section, we have seen that some DL-based NIDS models (Kitsune, ENIDrift) can be better interpreted in terms of decision trees than other models (HorusEye, HELAD). This might be attributed to the complexity of the NIDS model architecture and the properties of the model components. For example, HorusEye NIDS fidelity values are in the negative. Though HorusEye architecture includes only a single autoencoder in the second stage (which processes control plane traffic) compared to an ensemble of AEs in Kitsune or ENIDrift, that single AE is modified from a conventional AE to keep it lightweight, reduce the computational complexity, reduce the false-positive rate and improve throughput. It uses an asymmetric AE with separable convolution, dilation convolution and model quantization. As it is, AEs are limited in their interpretability as a result of their non-linear nature. Due to the complex architecture of the modified AE deployed in HorusEye, TRUSTEE might not be able to build a DT which can imitate the decisions of the black-box modified AE within an acceptable error rate. Therefore, to build DL-based NIDSs with better “explainability”, we recommend using less complex DL architectures and conventional components. However, there will be a trade-off between “explainability” and model performance (e.g., false-positive rate) and NIDS designers need to achieve a balance between the two.
@@
We have also seen from the results that TRUSTEE and SHAP explanations are in conflict for 2 out of 3 DL-based NIDSs considered in our work. Their explanations are in agreement for HorusEye NIDS but in conflict for Kitsune and ENIDrift NIDSs. This is probably due to the fact that SHAP value for each feature is calculated for a subset of data samples and reflects the importance of that feature in the decision made by the black-box model for a majority of the data samples in that subset. On the other hand, TRUSTEE acts on the complete input dataset and identifies the top features contributing the most to forming the decision boundary separating all benign and malicious data samples which are part of the input dataset. Another reason could be the different optimization criteria used by TRUSTEE and SHAP while building explanations. TRUSTEE applies a teacher-student dynamic derived from imitation learning that uses the black-box ML/DL model as an oracle in conjunction with a carefully curated dataset to guide the training of a surrogate “white-box” model in the form of a DT that imitates the black-box’s decisions. On the other hand, SHAP uses the Shapley kernel method to approximate the black-box model-interpretable model fidelity function which recovers Shapley values. The specific data subset used to conduct the TRUSTEE-SHAP comparison could also play a role in the conflict. To reduce the conflict between TRUSTEE and SHAP explanations at the subset-level, we need to further investigate its underlying cause in the future.
@@
Finally, it can be inferred from the results that some DL-based NIDSs (Kitsune) are more vulnerable to variable packet rate samples than others (HorusEye, HELAD) while for the rest (ENIDrift), such vulnerability remains inconclusive. One of the possible underlying reasons is that sometimes NIDSs are trained with training data containing higher than real-world attack traffic volume. To reduce vulnerability to variable packet rate samples, we recommend that NIDS designers use training data which reflects the attack traffic volume in real-world deployments. Additionally, the NIDS architecture may also have a role to play in its vulnerability to variable packet rate samples. An AE trained on benign traffic may be vulnerable to attack traffic with variable packet rates as the threshold used for output loss function (RMSE) is determined based on “in-distribution” attack traffic present in the training data. Since, Kitsune consists of an ensemble of AEs, the vulnerability effect is amplified. HorusEye consists of a single modified AE and therefore, its vulnerability to variable packet rate samples is not as pronounced as Kitsune. ENIDrift uses a different method of feature extraction (vector embedding) and updates the weights assigned to AEs in the ensemble based on change in attack traffic distribution which might keep it from producing a discernable pattern of dependence on attack traffic volume compared to Kitsune. We have observed earlier that HELAD NIDS anomaly scores are not vulnerable to variable packet rate samples to a significant extent. This may be attributed to HELAD’s architecture which combines an AE with an LSTM network. Since the final anomaly score is a weighted exponential aggregation of AE’s actual RMSE score and LSTM’s predicted RMSE scores for future samples (using historical samples). Even if the AE is vulnerable to variable packet rate samples, the LSTM’s predicted RMSE scores might be much larger compared to to the AE’s RMSE score (the exponential operation makes the value even larger), thus affecting the final anomaly score to a greater extent and reducing HELAD’s vulnerability to variable packet rate samples.
@@ VII Conclusion
We have analyzed four state-of-the-art DL-based NIDS models (Kitsune, HorusEye, ENIDRift, HELAD) using explainable AI techniques such as TRUSTEE and SHAP. Using both global explanations and local explanations for the blackbox models’ decisions, we have presented the most prominent features used by each NIDS model considered. We have compared the explanations generated across xAI methods for a given NIDS model (using our devised data subset-level explanation approach) as well as the explanations generated across the NIDS models for a given xAI method (using fidelity and other standard evaluation metrics). Finally, we evaluate the vulnerability of each NIDS model to inductive bias (variable packet rate samples). The results show that: (1) some DL-based NIDS models can be better interpreted in terms of decision trees than other models, (2) TRUSTEE and SHAP explanations are in conflict for most of the NIDS models considered in this work and (3) some NIDS models are more vulnerable to variable packet rate samples than other models.
@@ Appendix A More Experimental Results
In this section, we present explanations of the four state-of-the-art DL-based NIDSs (Kitsune, HorusEeye, ENIDrift, HELAD) generated using xAI methods such as TRUSTEE and kernel SHAP where the NIDSs were trained/tested with the CICIDS-2017 dataset.
@@ A-A Case Study- Kitsune $\times$ CICIDS-2017
TRUSTEE Analysis: The results of TRUSTEE analysis for Kitsune are shown in Table VII. Using TRUSTEE with $30\%$ of the original CICIDS-2017 dataset samples and no pruning results in the DT explanation which achieves 0.843 fidelity compared to Kitsune. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=3$) results in a DT explanation that is shown in Fig. 13 and achieves 0.919 fidelity compared to Kitsune.
@@
TABLE VII: TRUSTEE analysis results for Kitsune CICIDS-2017 model
@@
Detection of Inductive Bias: The CICIDS-2017 dataset does not provide a separate dataset for botnet traffic trace which could be used for comparison with the Mirai trace in terms of the attack traffic volume. However, using a similar approach for tampering as in the case of Mirai trace, we evaluated the following threshold ranges (in packets per second) for ARP requests originating from infected hosts : (1) 10-50, (2) 30-70 and (3) 50-90. The original trace as well as the tampered traces are shown in Fig. 15. It can be observed that: (1) there is no significant difference between the benign and attack traffic volumes for the original trace and (2) tampering had negligible effect on the difference between benign and attack traffic in terms of number of packets per second. Thus, unlike the Mirai trace, there is no significant difference between the benign and attack traffic volumes which can be learnt by the NIDS model and leave it vulnerable to a change in attack traffic distribution. Thus, we do not investigate the presence of inductive bias (vulnerability to variable packet rate samples) any further for NIDS models trained with the CICIDS-2017 dataset.
@@ A-B Case Study- HorusEye $\times$ CICIDS-2017
TRUSTEE Analysis: The results of TRUSTEE analysis for HorusEye are shown in Table VIII. Using TRUSTEE with $30\%$ of the CICIDS-2017 dataset samples with no pruning results in a DT explanation which achieves negative fidelity ($\sim-0.956$) compared to original HorusEye model. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation which achieves negative fidelity ($\sim-0.018$) compared to original HorusEye model but higher fidelity compared to no pruning. Negative fidelity values mean that the DT explanation fits HorusEye predictions worse than the mean of the predictions.
@@
TABLE VIII: TRUSTEE analysis results for HorusEye CICIDS-2017 model
@@ A-C Case Study- ENIDrift $\times$ CICIDS-2017
TRUSTEE Analysis: The results of TRUSTEE analysis for ENIDrift are shown in Table IX. Using TRUSTEE with $30\%$ of the CICIDS-2017 dataset samples and no pruning results in the DT explanation which achieves 0.929 fidelity compared to ENIDrift. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation (shown in Fig. 18) which achieves 0.53 fidelity compared to ENIDrift. This is expected since pruning the branches of a DT reduces its fidelity. The most prominent features ENIDrift uses to determine anomaly are: $f38$, $f81$, $f158$.
@@
TABLE IX: TRUSTEE analysis results for ENIDrift CICIDS-2017 model
@@ A-D Case Study- HELAD $\times$ CICIDS-2017
TRUSTEE Analysis: The results of TRUSTEE analysis for HELAD are shown in Table X. Using TRUSTEE with $30\%$ of the reduced CICIDS-2017 dataset samples and no pruning results in the DT explanation which achieves 0.0 fidelity compared to HELAD. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation (shown in Fig. 20) which achieves 0.0 fidelity compared to HELAD.
@@
TABLE X: TRUSTEE analysis results for HELAD NIDS